waths the solution to 3x + 12=8x - 8
Found 2 solutions by superkid208, Nate:
Answer by superkid208(3)

3x+12=8x-8
3x+12-12=8x-8-12
3x=8x-20
3x-8x=8x-8x-20
-5x=-20
-5x/5=-20/5
x=-4

Answer by Nate(3500)

3x + 12 = 8x - 8
20 = 5x
4 = x